OpenStudy (anonymous):

I need help on this question please :) Suppose that x = 6 and y = 8. If y is inversely related to x, what is y when x = 3? a.)4 b.)5 c.)16 d.)18

OpenStudy (anonymous):

@him1618

OpenStudy (anonymous):

y = k/x k = yx Find K = 6*8 = 48 y = k/x y = 48/3 y =?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

np